Jocelynnn:

In a race , Ram covers 5 km in 20 min. How much distance will he cover in 100 min ?

AZ:

5 km in 20 minutes ?? km in 100 minutes How much distance would Ram cover?

AZ:

We can set it up as a proportion \(\dfrac{5}{20} = \dfrac{x}{100}\) Or if you want, Ram runs 5 km in 20 minutes If he runs 100 minutes, that's 5 times longer right? 20 * 5 = 100 minutes So then that means he runs 5 times the distance So what is 5 * 5 = ??